OpenStudy (anonymous):

solve for X, show work 5x + 7 + 3x = -8 + 3x

OpenStudy (anonymous):

add the x's on either side of the equation and tell me what you get.

OpenStudy (anonymous):

oh you ._. heres what i got 5x + 3x - 3x + 7 = -8 8x - 3x + 7 = -8 5x = -8 - 7 5x = 15 x = 3

OpenStudy (anonymous):

Close; you swapped the signs on the -8 -7 to 15 not -15.

OpenStudy (anonymous):

meh. thanks guy

OpenStudy (anonymous):

word. good luck.

OpenStudy (anonymous):

12 sheets to have finished by 4 pm, i need all the luck i can get

OpenStudy (anonymous):

I'll be on for a while so I'll try to help if I can.

OpenStudy (anonymous):

haha, thanks